OpenStudy (anonymous):

The Tigers won 8 games more than they lost, and there were no ties. If the Tigers played 78 games, how many games did they loose?

jimthompson5910 (jim_thompson5910):

"The Tigers won 8 games more than they lost" ----> W = L + 8

jimthompson5910 (jim_thompson5910):

"the Tigers played 78 games (with no ties)" ----> W + L = 78

jimthompson5910 (jim_thompson5910):

W + L = 78 L + 8 + L = 78 Keep going to find L

OpenStudy (anonymous):

im so confused

jimthompson5910 (jim_thompson5910):

What do you get when you solve for L?

OpenStudy (anonymous):

i dont get what you did

OpenStudy (anonymous):

i wrote let w= the games won let l= the games lost w+l=79

OpenStudy (anonymous):

thats all i get...

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

They played 78 games, not 79

jimthompson5910 (jim_thompson5910):

So W + L = 78

jimthompson5910 (jim_thompson5910):

L + 8 + L = 78 2L + 8 = 78 2L + 8-8 = 78-8 2L = 70 what's next?

OpenStudy (anonymous):

where does the l+8+l+8 come from? thats the only thing i dont get

OpenStudy (anonymous):

L + 8 + L = 78

jimthompson5910 (jim_thompson5910):

I have W + L = 78

jimthompson5910 (jim_thompson5910):

Replace W with L + 8 (since W = L + 8) to get L + L + 8 = 78

OpenStudy (anonymous):

oh okay i got it so the anwer is 35 games lost

OpenStudy (anonymous):

thanks so much

jimthompson5910 (jim_thompson5910):

35 is the correct answer